OpenStudy (anonymous):

how do you change the following point into rectangular form (-2, 5pi/3) A. (-1, sqrt3) B. (-2, -sqrt3/2) C. (-2, sqrt3) D. none of the above

OpenStudy (ddcamp):

This is in polar (r,ø) form. To convert from polar to rectangular, use the formulas: \[x = r\cos(\theta) \\ y = r\sin(\theta)\]

OpenStudy (anonymous):

In the expression, (-2, 5pi/3) -2 is r 5pi/3 is theta you have formula to convert it to rectangular form, those are x = r cos theta y = r sin theta. apply and show me your work